OpenStudy (anonymous):

The student's response is wrong. What did the student do incorrectly? Question: Simplify by using order of operations 7(-9 + 10) - 3 + 4 Student's Answer: 7(1) - 3 + 4 = 7 - 3 + 4 = 7 - 7 = 0

OpenStudy (anonymous):

You have to do whats in the parenthesis first.

OpenStudy (anonymous):

Then you multiply, then you add/subtract from left to right.

OpenStudy (anonymous):

did she do anything wrong though?

OpenStudy (anonymous):

Yes she did. 7-3+4 = 4+4 = 8

OpenStudy (anonymous):

but i thought she was suppose to add first?

OpenStudy (anonymous):

oh nvm thanks

OpenStudy (anonymous):

no, subtraction and addition goes from left to right, only multiplication and division comes before, those two. You're welcomed :)